# Measurement and feedforward correction of the fast phase noise of lasers

### v1

@@ III.4 Gallery of laser frequency noise
As expected, the fast frequency noise of these lasers drastically depends on their architecture, with the DBR laser reaching a white frequency noise floor of 10 000 Hz2/Hz while the solid-state laser noise floor is below the sensibility of 0.01 Hz2/Hz of our detection scheme. The ECDL is around 100 Hz2/Hz, limited as the DBR by the fundamental quantum noise of lasers given by the Schawlow-Townes formula $S_{\nu}^{\rm ST}=h\nu_{0}\nu_{c}^{2}/P$, with $P$ the output power, $\nu_{0}$ the laser frequency, and $\nu_{c}$ the laser cavity linewidth [36]. This noise is especially high for ECDL and DBR due to the limited power that can be extracted from side-emitting diodes (below 100 mW) as well as by the large cavity bandwidth (GHz or more). In contrast, a VECSEL would have orders of magnitude smaller Schawlow-Townes noise-floor thanks to the multi-Watt power accessible with the surface-emitting design as well as by the MHz linewidth of their cavity (associated to the much lower gain in this design). In fact, the measured noise level of the VECSEL at 10 Hz2/Hz is probably limited by thermal and carrier-density fluctuation created by intensity noise of the pump, rather than by the Schawlow-Townes limit [37, 18]. For the TiSapph laser, its noise level is so low that we cannot measure it with our finite measurement noise floor, in agreement with its expected Schawlow-Townes level at $10^{-5}$ Hz2/Hz. The SIL laser display a noise level of 0.6 Hz2/Hz, which is explained by the self-stabilization mechanism of the laser when coupled to a high-Q resonator [38, 19, 39]. Unfortunately, no fiber-based laser was available at the time of these measurements, but one could expect frequency noise at a level below 10 Hz2/Hz.
@@
For all lasers, expect the VECSEL, we used home-made MZIs described previously and all measurements have in common a displayed noise level largely increasing below 100 kHz, originating from the MZI and not the lasers. In contrast, for the VECSEL, we used a commercial phase noise analyzer (OE4000 from OEwaves), whose performance remains excellent at lower Fourier frequencies. We note that, at MHz Fourier frequencies, the commercial system detection noise floor is 1 Hz2/Hz, while the simple MZI can reach a 0.1 Hz2/Hz level. We could even push down to 0.01 Hz2/Hz (see the TiSapph curve in Fig. 3(a)), by using a single biased photo-detector and were then limited by intensity noise of the laser. We also compare, in Fig. 3(e), the measured noise with our home-made MZI and with the commercial phase noise analyzer for the ECDL laser. The agreement is very good, except in the 10 kHz range where the MZI exhibits large noise, described earlier.
@@
In Fig. 3(b), we compare the noise spectra of four lasers of identical architecture (ECDLs), from the same manufacturer but with different wavelengths, and obtain very similar frequency noise characteristics. In Fig. 3(c), we compare the fast frequency noise of a laser before and after frequency doubling using SHG (second harmonic generation) in an enhancement cavity (with a linewidth of  100 kHz): as expected, the noise power is multiplied by 4 at low Fourier frequency, but is filtered above the cut-off frequency of the cavity. This filtering action of a narrow linewidth cavity is used in Ref. [13] to passively reduce the noise of diode lasers. In Fig. 3(d), we illustrate the effect of a feedback loop on the noise reduction within the loop bandwidth: for the fastest (PI corner at 500 kHz) and highest gain setting, the noise can be reduced below 300 kHz, but a servo bump appears at 1 MHz. In the next section, we will see how the fast feedforward noise eater is effective in this region where the feedback approach fails.
@@ IV.1 Theory
The range of Fourier frequency $30f_{c}<f<0.15/\tau$ of efficient feedforward operation can be widened by decreasing the cut-off frequency $f_{c}$ and the MZI delay $\tau$, however this comes at the cost of decreasing the MZI sensibility $s_{0}$ and more filtering of its output by the LP filter, such that a larger gain $G$ is required.
@@ IV.3 Results
Figure 4: Feedforward correction results. (a) Elements in the feedforward path. (b) Noise correction (in ratio of PSD) achieved experimentally (light blue) compared to the estimated limits set by the noise-floor (dark blue), LP filter dispersion (dot-dashed line), MZI discriminator roll-off (dashed), delay error (dotted) and the sum of these (solid black line) (c) Estimated noise reduction $|1-C|^{2}$ (in dB) as a function of Fourier-frequency $f$ (top) originating from the phase error $\arg{C}$ (middle) and gain error $|C|$ (bottom). Three different length of delay line are shown. (d) Noise reduction for several LP cut-off frequencies $f_{c}$ and model (dot-dashed). (e) Noise reduction for several delays and model (dotted).
@@
We now go into each of the contributions to the full model. First, the finite noise floor of the detection, originating from the photo-diode transimpendance amplifier, sets a frequency-independent limit of the correction to 33 dB (dark blue measurement) which dominates around $f=3-5$ MHz. Second, at lower frequency, the correction is limited by the LP filter dispersion (Fourier-frequency-dependent phase-lag), as demonstrated by the perfect match with the dot-dashed line. We further illustrate this by varying the cut-off frequency ($f_{c}$) in Fig. 4(d). Third, at the highest frequencies, the MZI response roll-off explains the degrading performance of the feedforward correction (dashed line). Fourth, we show with the dotted line the contribution of a 0.5 ns mismatch of delay (the smallest resolution we had), which is not expected to contribute at this level. For illustration, we also show, in Fig. 4(e), how the correction is affected by larger delays of 4 and 8 nanoseconds set on purpose. Finally, by including all four independent imperfections into a single curve (solid line), we obtain a good agreement with the measured correction ratio.
@@
Following these observations, further improvements of the phase noise eater can be envisioned. The low and high frequency response can be improved by having a slower filter (decrease $f_{c}$) and faster MZI discriminator (decrease $\tau$), and by optimizing the amplifier to compensate for the decreased sensibility. Also, an automatic gain control procedure acting on the continuously-variable optical attenuator, would allow to keep optimized conditions even for drifting laser power at the MZI input that otherwise disturbs the overall gain [40].
@@
Finally, we comment on the influence of the phase noise eater on intensity noise. Electro-optic modulators (either free-space or waveguide-based) used for applying the phase correction also slightly affect the intensity, an effect called Residual Amplitude Modulation (RAM). For our device, we measured a ratio of intensity to phase modulation of -59 dB (changing the phase by 1 rad affects the intensity by 0.1 %). We note that the RAM depends strongly on the DC bias applied to the EOM (probably caused by some etalon effect), and can be further decreased by 30 dB for an optimal bias. Operating at such a sweet spot would however require a dedicated locking procedure of the bias point as it otherwise drifts. For this discussion, we picked the point of maximum RAM. The excess RIN (relative intensity noise) caused by the feedforward correction is simply ${\rm RIN}={\rm RAM}\times S_{\phi}$. In practice, for the laser shown in Fig. 1, with a measured frequency noise of 1000 Hz2/Hz and RIN of -150 dBc/Hz, we observed that the RIN doubles at $f=1$ MHz (the excess RIN from the feedforward is equal to the laser RIN). We conclude by pointing out that such low level of RIN have a negligible contribution to the error budget of quantum gates [11].
@@ IV.4 Discussion
Finally, other feedforward noise suppression results have already been reported in the litterature. With the same configuration as reported here, Hashemi and coworkers reduced the frequency noise of a DBR laser with a white noise floor at $S_{\nu}\sim 10^{6}\,{\rm Hz}^{2}/{\rm Hz}$ by 15 dB for $f$ up to 100 MHz [29, 30]. This large bandwidth was made possible by using a smaller delay in the interferometer at the cost of a reduced sensitivity (and thus increased noise floor), which is reasonable there given the large initial noise of the DBR laser used in their work. Other approaches than using a delayed interferometer for frequency noise measurement and feedforward are also possible. In Ref. [32], the authors perform an heterodyne measurement between the noisy laser and its filtered version at the output of a cavity, and demonstrated 20 dB suppression at sub-MHz Fourier frequencies and 3 dB at 5 MHz. In Refs. [33, 40], Chao, Hua and coworkers instead directly use the same Pound-Drever-Hall error signal as used for feedback stabilization on a reference cavity, and feedforward its high-frequency components, directly proportional to phase noise, to an electro-optic modulator. They achieve 30 dB of cancellation in the $0.3-3$ MHz range, a broader range than with our MZI measurement thanks to the wider region where the PDH signal is proportional to phase noise (for $f_{c}\ll f\ll{\rm FSR}$). The absolute noise was however not reported (only the relative cancellation for an intentionally added modulation), neither the measurement noise floor, so it is yet unclear if the noise can be pushed below the 10 Hz2/Hz level. Such cancellation performance were also later reproduced independently, achieving 30 dB at 2 MHz [41].
@@ VI Conclusion and Outlook
In the second part of this work, we showed how the phase noise measurement can be used for a feedforward correction of the phase. We achieved up to 30 dB of cancellation at a Fourier frequency of 3 MHz, and more than 20 dB in the 1-10 MHz, bringing the laser phase noise below the Hz2/Hz level. The performance of the system was discussed and matched well with calculations. We finally highlighted the effect of improving the laser phase noise when manipulating the coherence of single neutral atoms.

### v2

@@ III.4 Gallery of laser frequency noise
As expected, the fast frequency noise of these lasers drastically depends on their architecture, with the DBR laser reaching a white frequency noise floor of 10 000 Hz2/Hz while the solid-state laser noise floor is below the sensitivity of 0.01 Hz2/Hz of our detection scheme. The ECDL is around 100 Hz2/Hz, limited as the DBR by the fundamental quantum noise of lasers given by the Schawlow-Townes limit: $S_{\nu}^{\rm qu}=h\nu_{0}\nu_{c}^{2}/P\times(1+\alpha^{2})$, with $P$ the output power, $\nu_{0}$ the laser frequency, $\nu_{c}$ the laser cavity linewidth; and also including the Henry enhancement factor $\alpha$ (typically between 2 to 7 for semiconductor lasers and lower for solid state lasers) [36]. This noise is especially high for ECDL and DBR due to the limited power that can be extracted from side-emitting diodes (below 100 mW) as well as by the large cavity bandwidth (GHz or more). In contrast, a VECSEL would have orders of magnitude smaller Schawlow-Townes noise-floor thanks to the multi-Watt power accessible with the surface-emitting design as well as by the MHz linewidth of their cavity (associated to the much lower gain in this design). In fact, the measured noise level of the VECSEL at 10 Hz2/Hz is probably limited by thermal and carrier-density fluctuation created by intensity noise of the pump, rather than by the Schawlow-Townes limit [37, 18]. For the TiSapph laser, its noise level is so low that we cannot measure it with our finite measurement noise floor, in agreement with its expected Schawlow-Townes level at $10^{-5}$ Hz2/Hz. The SIL laser display a noise level of 1 Hz2/Hz, which is explained by the self-stabilization mechanism of the laser when coupled to a high-Q resonator [38, 19, 39]. Unfortunately, no fiber-based laser was available at the time of these measurements, but one could expect frequency noise at a level below 10 Hz2/Hz [40].
@@
For all lasers, expect the VECSEL, we used home-made MZIs described previously and all measurements have in common a displayed noise level increasing below 100 kHz, originating from the MZI and not the lasers. In contrast, for the VECSEL, we used a commercial phase noise analyzer (OE4000 from OEwaves), whose performance remains excellent at lower Fourier frequencies. We note that, at MHz Fourier frequencies, the commercial system detection noise floor is 1 Hz2/Hz, while the simple MZI can reach a 0.1 Hz2/Hz level. We could even push down to 0.01 Hz2/Hz (see the TiSapph curve in Fig. 3(a)), by using a single biased photo-detector and were then limited by photon shot noise. We also compare, in Fig. 3(e), the measured noise with our home-made MZI and with the commercial phase noise analyzer for the SIL laser. The agreement is very good, except in the 10 kHz range where the MZI exhibits an excess noise, described earlier.
@@
In Fig. 3(b), we compare the noise spectra of four lasers of identical architecture (ECDLs), from the same manufacturer but with different wavelengths, and obtain very similar frequency noise characteristics. In Fig. 3(c), we compare the fast frequency noise of an ECDL laser before and after frequency doubling from 960 to 480 nm using SHG (second harmonic generation) in an enhancement cavity (with a linewidth of  100 kHz): as expected, the noise power is multiplied by 4 to 10 at low Fourier frequency, but is filtered above the cut-off frequency of the cavity. This filtering action of a narrow linewidth cavity is used in Ref. [13] to passively reduce the noise of diode lasers. In Fig. 3(d), we illustrate the effect of a feedback loop on the noise of a DBR laser: for the fastest (PI corner at 500 kHz) and highest gain setting, the noise can be reduced below 300 kHz, but a servo bump appears at 1 MHz. In the next section, we will see how the fast feedforward noise eater is effective in this region where the feedback approach fails.
@@ IV.1 Theory
The range of Fourier frequency $30f_{c}<f<0.15/\tau$ of efficient feedforward operation can be widened by decreasing the cut-off frequency $f_{c}$ and the MZI delay $\tau$, however this comes at the cost of decreasing the MZI sensitivity $s_{0}$ and more filtering of its output by the LP filter, such that a larger gain $G$ is required.
@@ IV.3 Results
Figure 4: Feedforward correction results. (a) Elements in the feedforward path. (b) Noise correction (in ratio of PSD) achieved experimentally (light blue) compared to the estimated limits set by the noise floor (dark blue), LP filter dispersion (dot-dashed line), MZI discriminator roll-off (dashed), delay error (dotted) and the sum of these (solid black line). (c) Estimated noise reduction $|1-C|^{2}$ (in dB) as a function of Fourier-frequency $f$ (top) originating from the phase error $\arg{C}$ (middle) and gain error $|C|$ (bottom). Three different length of delay line are shown. (d) Noise reduction for several LP cut-off frequencies $f_{c}$ and model (dashed). (e) Noise reduction for several delays and model (dotted).
@@
We now go into each of the contributions to the full model. First, a frequency-independent limit of correction of 32 dB (dark blue line) is set by the photon shot noise and the electronic noise floor of the photo-detector that are both on the order $\rm 0.1~{}Hz^{2}/Hz$ (see Fig. 2). This limit is reached around $f=3-5$ MHz. Second, at lower frequency, the correction is limited by the LP filter dispersion (Fourier-frequency-dependent phase-lag), as demonstrated by the perfect match with the dot-dashed line. We further illustrate this by varying the cut-off frequency ($f_{c}$) in Fig. 4(d). Third, at the highest frequencies, the MZI response roll-off explains the degrading performance of the feedforward correction (dashed line). Fourth, we show with the dotted line the contribution of a 0.5 ns mismatch of delay (the smallest resolution we had), which is not expected to contribute at this level. For illustration, we also show, in Fig. 4(e), how the correction is affected by larger delays of 4 and 8 nanoseconds set on purpose. Finally, by including all four independent imperfections into a single curve (solid line), we obtain a good agreement with the measured correction ratio.
Having discussed the improved phase stability in the 0.1-10 MHz range, we now consider if significant noise could have been added in the lower frequency range. One contribution could be the noise of the first MZI which is then fed to the fiber-EOM. As seen in Fig. 3(e), the MZI noise is at the $10^{2}$ Hz2/Hz ($10^{4}$ Hz2/Hz) level at $f=10$ kHz (1 kHz). If this noise would be imprinted directly to the laser, it would contribute a gate error of $3\times 10^{-5}$ [11]. However, we realize the integrator performing frequency-to-phase conversion before the phase-modulator (fiber-EOM) with a low-pass filter of cut-off frequency $50$ kHz, which is equivalent to an ideal integrator followed by a high-pass filter. Consequently, the MZI noise imprinted onto the laser is largely suppressed below 10 kHz, further reducing its contribution to gate error by a factor of more than 30, making this noise source negligible in the current noise budget [11]. A second contribution could be the noise imprinted by the long fiber delay line (10 m), which is sensitive to thermo-mechanical noise. Previous studies have measured excess frequency noise on a level of $10^{2}$ Hz2/Hz in presence of 80 dB of acoustic noise [42, 43]. While such noise would be deleterious when operating at low Rabi frequencies (such as in metrology), for gate operations with MHz-range Rabi frequencies it is again negligible.
@@
Following these observations, further improvements of the phase noise eater can be envisioned. The low and high frequency response can be improved by having a slower filter (decrease $f_{c}$) and faster MZI discriminator (decrease $\tau$), and by optimizing the amplifier to compensate for the decreased sensitivity . Also, an automatic gain control procedure acting on the continuously-variable optical attenuator, would allow to keep optimized conditions even for drifting laser power at the MZI input that otherwise disturbs the overall gain [41].
@@
Finally, we comment on the influence of the phase noise eater on intensity noise. Electro-optic modulators (either free-space or waveguide-based) used for applying the phase correction also slightly affect the intensity, an effect called Residual Amplitude Modulation (RAM). For our device, we measured a ratio of phase to intensity modulation of $k=\rm 0.1\%/rad$. We note that the RAM depends strongly on the DC bias applied to the EOM (probably caused by some etalon effect), and can be further decreased by a factor 30 for an optimal bias. Operating at such a sweet spot would however require a dedicated locking procedure of the bias point as it otherwise drifts. For this discussion, we picked the point of maximum RAM. The excess RIN (relative intensity noise) caused by the feedforward correction is simply ${\rm RIN}=k^{2}\times S_{\phi}$. In practice, for the laser shown in Fig. 1, with a measured frequency noise of 1000 Hz2/Hz and RIN of $-150$ dBc/Hz, we observed that the RIN doubles at $f=1$ MHz (the excess RIN from the feedforward is equal to the laser RIN). We conclude by pointing out that such low level of RIN have a negligible contribution to the error budget of quantum gates [11].
@@ IV.4 Discussion
Finally, other feedforward noise suppression results have already been reported in the literature. With the same configuration as reported here, Hashemi and coworkers reduced the frequency noise of a DBR laser with a white noise floor at $S_{\nu}\sim 10^{6}\,{\rm Hz}^{2}/{\rm Hz}$ by 15 dB for $f$ up to 100 MHz [29, 30]. This large bandwidth was made possible by using a smaller delay in the interferometer at the cost of a reduced sensitivity (and thus increased noise floor), which is reasonable there given the large initial noise of the DBR laser used in their work. Other approaches than using a delayed interferometer for frequency noise measurement and feedforward are also possible. In Ref. [32], the authors perform an heterodyne measurement between the noisy laser and its filtered version at the output of a cavity, and demonstrated 20 dB suppression at sub-MHz Fourier frequencies and 3 dB at 5 MHz. In Refs. [33, 41], Chao, Hua and coworkers instead directly use the same Pound-Drever-Hall error signal as used for feedback stabilization on a reference cavity, and feedforward its high-frequency components, directly proportional to phase noise, to an electro-optic modulator. They achieve 30 dB of cancellation in the $0.3-3$ MHz range, a broader range than with our MZI measurement thanks to the wider region where the PDH signal is proportional to phase noise (for $f_{c}\ll f\ll{\rm FSR}$). The absolute noise was however not reported (only the relative cancellation for an intentionally added modulation), neither the measurement noise floor, so it is yet unclear if the noise can be pushed below the 10 Hz2/Hz level. Such cancellation performance were also later reproduced independently, achieving 30 dB at 2 MHz [44].
@@ VI Conclusion and Outlook
In the second part of this work, we showed how the phase noise measurement can be used for a feedforward correction of the phase. We achieved up to 30 dB of cancellation at a Fourier frequency of 3 MHz, and more than 20 dB in the 1-10 MHz, bringing the laser phase noise below the Hz2/Hz level. The performance of the system was discussed and matched well with calculations. We finally highlighted the effect of improving the laser phase noise when manipulating the coherence of an array of individual neutral atoms.